OpenStudy (anonymous):

how to solve the following 1) cos( sin^ -1 (-3/5) ) 2)cosec (cos^-1(-12/13))

OpenStudy (anonymous):

@mathmate @ganeshie8 @Loser66 @e.mccormick @eyust707 @Preetha @wolfe8 @iPwnBunnies @paki @ShadowLegendX

OpenStudy (loser66):

what is it? it's not an equation to solve. At least, you have to have = sign on the expression

OpenStudy (anonymous):

@Loser66 they are two different sums , we have to find the value of

OpenStudy (loser66):

I am sorry for my stupidness, I don't understand still. @Hero

hero (hero):

\(\cos\left((\sin^{-1}\left(-\dfrac{3}{5}\right)\right)\) right?

OpenStudy (anonymous):

yes right @Hero

hero (hero):

First we have to find \(\theta\) since \(\sin(\theta) = -\dfrac{3}{5}\)

hero (hero):

Once you find theta, then you have to find \(\cos(\theta)\)

OpenStudy (anonymous):

so what will be θ @Hero

hero (hero):

If you can't use a calc, just set it up this way: